OpenStudy (anonymous):

does 2thesymbolforpieR mean 2 times pie times r

OpenStudy (wolf1728):

The symbol for pi is just this π It means 3.14159265 that is all - no 2 multiplied in there.

OpenStudy (wolf1728):

I don't know why you typed 2thesymbolforpieR Anyway, pi is just π

OpenStudy (wolf1728):

Maybe you were trying to type something like this? 2 PI r which is 2 times pi times the radius of the circle which equals the circumference.

OpenStudy (anonymous):

I understand although the question says find the area of a circle with a given diameter of 12.56. my computer doesnt have a pie symbol.

OpenStudy (anonymous):

so it will be 2 times pie times 12.56?

OpenStudy (wolf1728):

no because cirumfeence = PI * diamter

OpenStudy (anonymous):

and the question is in diameters.

OpenStudy (wolf1728):

well circumference is strictly pi*diameter

OpenStudy (wolf1728):

by the way to make the pi symbol hold down the alt key and at the same time type 227 and you will get π

OpenStudy (anonymous):

I'm not trying to find circumference, I'm trying to find the area. Please explain?

OpenStudy (wolf1728):

Can you make the pi symbol yet? π Anyway circle area = π • r²

OpenStudy (anonymous):

I tried it, it's not working

OpenStudy (wolf1728):

You know the question and the area?

OpenStudy (wolf1728):

or did you mean the π symbol

OpenStudy (anonymous):

I meant the pie symbol.

OpenStudy (wolf1728):

Did you use the keypad on the right side of the keyboard? If you use the numbers at the top of the keyboard, it won't work.

OpenStudy (anonymous):

▬0♠♣♦☻☺◘○I'm trying to find the area in diameters.

OpenStudy (anonymous):

←π

OpenStudy (anonymous):

yahoo I got it!!!!

OpenStudy (anonymous):

now back to the question

OpenStudy (wolf1728):

You don't find the area in diameters. Anyway, diameter = 12.56 so radius = 12.56 / 2 = 6.28 Area would then equal 3.14159 * 6.28 * 6.28

OpenStudy (anonymous):

I'm sry ur dividing, daaaaaaaaaaaaa

OpenStudy (anonymous):

why do u multiply it by pie twice

OpenStudy (wolf1728):

I am not multiplying it by anything diameter = 12.56 radius = one half diameter = 12.56 / 2 = 6.28 area of a circle is π•r² so area = π • 6.28 • 6.28 = 123.9

OpenStudy (anonymous):

;thanks Im gonna have to make flash cards.

OpenStudy (wolf1728):

okay - glad to help you out you'll find circle area is easy as π

OpenStudy (anonymous):

funny

OpenStudy (wolf1728):

thanks :-)

OpenStudy (anonymous):

Why does my book have c=2πr? Is it a given that pie always stays pie?

OpenStudy (anonymous):

to find the area, I take πr to the second power?

OpenStudy (anonymous):

I got 6590, well actually 6589.930414.

OpenStudy (wolf1728):

circumference = 2 • π • r area = π • r²

OpenStudy (wolf1728):

You do not tkae π•r to the second power only 'r' to the second power

OpenStudy (wolf1728):

π • r to the second power is π •π • r • r but that is NOT how circle area is calculated It is calculated by π • r • r

OpenStudy (wolf1728):

Well I came back here. I hope you note what I posted.

OpenStudy (anonymous):

I did π * 6.28 * 6.28=123.9 ft.?

OpenStudy (wolf1728):

Yes that's what it equals. Actually it would be 123.9 square feet